# Globally-Constrained Decentralized Optimization with Variable Coupling

### v3

@@ Abstract
Many realistic decision-making problems in networked scenarios, such as formation control and collaborative task offloading, often involve complicatedly entangled local decisions, which, however, have not been sufficiently investigated yet. Motivated by this, we study a class of decentralized optimization problems with a variable coupling structure that is new to the literature. Specifically, we consider a network of nodes collaborating to minimize a global objective subject to a collection of global inequality and equality constraints, which are formed by the local objective and constraint functions of the nodes. On top of that, we allow such local functions of each node to depend on not only its own decision variable but the decisions of its neighbors as well. To address this problem, we propose a decentralized projected primal-dual algorithm. It first incorporates a virtual-queue technique with a primal-dual-primal scheme, and then linearizes the non-separable objective and constraint functions to enable decentralized implementation. Under mild conditions, we derive $O(1/k)$ convergence rates for both objective error and constraint violations. Finally, two numerical experiments corroborate our theoretical results and illustrate the competitive performance of the proposed algorithm.
Index Terms:
Decentralized optimization, globally-coupled constraint, primal-dual algorithm, variable coupling.
@@ I Introduction
Decentralized optimization has garnered considerable recent attention due to its broad applications in various networked systems such as communication networks[1], smart grids[2], and computing networks[3]. In these networked systems, each node possesses a collection of local data, typically inaccessible to other nodes for privacy protection, and all the nodes aim to make optimal decisions by solving a global optimization problem determined by all their local data. Decentralized optimization techniques allow the nodes to collaboratively solve such a problem through communicating with their neighboring nodes only, leading to high scalability with respect to the network size and the data volume.
@@
To date, a host of decentralized optimization algorithms have been developed. Among these works, the majority focus on consensus optimization, which aims at finding a common decision that minimizes the sum of the local objectives under, mostly, no constraints [4, 5] or rather simple constraints such as set constraints [6, 7], local equality/inequality constraints[8, 9], and equality/inequality constraints known to all the nodes [10, 11]. Another line of research explores problems with local variables subject to coupling equality/inequality constraints [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. In particular, to solve the challenging problems with globally-coupled nonlinear inequality constraints consisting of all the local constraint functions from the nodes, the distributed primal-dual subgradient algorithms [15, 16, 23, 22, 20, 24, 19], the distributed method based on operator splitting[18], the distributed integrated primal-dual proximal algorithm [25], and the relaxation and successive distributed decomposition method[17] along with its extension [21] are developed. These works also derive various convergence results, including asymptotic convergence [15, 16, 17, 18, 19] as well as $O(\ln k/\sqrt{k})$ [20], $O(1/\ln k)$ [21], $O(1/\sqrt{k})$ [22, 23, 24], and $O(1/k)$ [25] convergence rates.
@@
A common limitation of the aforementioned decentralized optimization methods [4, 5, 6, 7, 12, 13, 8, 9, 10, 11, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] is that each node’s local objective and constraint functions only rely on its own local decision variable. This excludes many realistic problems where each node’s local function values are influenced by the decisions made by other nodes in the network, especially the node’s neighbors. A representative example is the task offloading problem over cooperative mobile edge computing networks [26], where each edge server’s task completion time and local resource consumption are affected by the task offloaded from the neighboring servers. Such a variable coupling structure also arises in collaborative resource allocation [27], distributed model predictive control [28], wireless localization [29], etc.
@@
Motivated by this, we consider a decentralized optimization problem with globally-coupled nonlinear inequality and linear equality constraints. Moreover, for each node, we allow both the local objective function and the local constraint functions to involve variable coupling, i.e., these local functions are associated with the decision variables of the node itself and all its neighbors. This coupling structure endows the optimization problem with a relatively general form that generalizes the prior prevalent decentralized optimization models in [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], which do not admit variable coupling in the local functions, and the formulations of the application problems in [26, 27, 28], which admit variable coupling but involve no globally-coupled constraints. A few existing works [30, 31, 32, 33, 34, 35] also consider variable coupling, but they either focus on simpler problems (e.g., [30] on unconstrained problems and [31] on convex feasibility problems) or different variable coupling models [32, 33, 34, 35] that cannot generalize ours (cf. Section II).
@@
In this paper, we develop a decentralized projected primal-dual algorithm for addressing the aforementioned globally-constrained optimization problem with variable coupling. To this end, we first construct a variant of the augmented Lagrangian function with respect to an equivalent form of the original problem, whose gradient has a certain degree of separability. Then, the primal update is designed so as to cause a descent of the augmented Lagrangian variant, where the gradient projection operation is employed. The dual updates can be viewed as a non-trivial integration of ideas from the virtual queue method [36] and a dual variant of the P-EXTRA algorithm [37]. Eventually, the variable coupling structure and the globally-coupled constraints are effectively decoupled, so that our proposed algorithm can be executed in a fully decentralized fashion. We show that the proposed algorithm achieves $O(1/k)$ convergence rates in terms of both objective error and constraint violations under mild conditions. Its effectiveness and efficiency are demonstrated by two numerical examples.
@@
  • Compared to the existing works [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] that are able to solve problems with global inequality constraints but no variable coupling (i.e., special cases of our problem), the $O(1/k)$ convergence rate of our proposed algorithm is stronger than the asymptotic convergence established in [15, 16, 17, 18, 19], faster than the convergence rates in [20, 22, 21, 23, 24], and comparable to the rate in [25].
  • Our algorithm has lower computational costs per iteration than the algorithms in [25, 24, 20, 19, 21, 17, 16, 22] because it does not need to solve an optimization subproblem per iteration. Additionally, our algorithm adopts constant step-sizes, which usually leads to faster convergence in practice compared to the algorithms with diminishing step-sizes [15, 16, 17, 20, 21, 22, 23, 24].
@@
[…]
@@ II Problem Formulation
For each $i\in\mathcal{V}$, $x_{i}\in\mathbb{R}^{d_{i}}$ is the decision variable of node $i$ and is restricted to node $i$’s constraint set $X_{i}\subseteq\mathbb{R}^{d_{i}}$. Additionally, $f_{i}:\mathbb{R}^{\sum_{j\in\mathcal{N}_{i}}d_{j}}\rightarrow\mathbb{R}$ is the local objective function and $g_{i}:\mathbb{R}^{\sum_{j\in\mathcal{N}_{i}}d_{j}}\rightarrow\mathbb{R}^{p}$ is the local inequality constraint function of node $i$, which are known only to node $i$ but are associated with the decision variables of both node $i$ and its neighbors, denoted by $x_{\mathcal{N}_{i}}$. Here, each $g_{i}$ is a vector-valued function, and we express it as $g_{i}\left(x_{\mathcal{N}_{i}}\right)=[g_{i1}\left(x_{\mathcal{N}_{i}}\right),\ldots,g_{ip}\left(x_{\mathcal{N}_{i}}\right)]^{T}$. Moreover, $A_{i}\in\mathbb{R}^{m\times\sum_{j\in\mathcal{N}_{i}}d_{j}}$ and $b_{i}\in\mathbb{R}^{m}$, privately owned by node $i$, encode node $i$’s local component of the global equality constraint in (1), which, again, involves the decision variables of all the nodes in $\mathcal{N}_{i}$.
@@
Compared to most distributed optimization models[4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 30, 31, 32, 33, 34, 35] in the literature, problem (1) has a more complicated and general coupling structure and, thus, is more challenging to solve. Specifically, the problems considered in many existing works [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] are “separable” in the sense that each local function belonging to node $i$ is only associated with node $i$’s decision variable $x_{i}$. In contrast, problem (1) is non-separable, as each $f_{i}$, $g_{i}$, and $A_{i}$ are associated with $x_{\mathcal{N}_{i}}$, i.e., $x_{j}$ $\forall j\in\mathcal{N}_{i}$. In addition, $g_{i}$, $A_{i}$, $b_{i}$ $\forall i\in\mathcal{V}$ constitute the globally-coupled inequality and equality constraints in an additive manner. Such a non-separable structure is not uncommon in networked applications, as each node’s cost, utility, capacity, etc., are often influenced by its neighbors’ states and play a role in the network-wide problems. Similar variable coupling structures are also considered in [30, 31], but [30] only investigates unconstrained convex optimization and [31] studies a convex feasibility problem.
@@
There have been several other coupling models for distributed convex optimization with an additive global objective function. In [32] and [33], each element in a global decision vector is assigned to a subset of nodes that are not necessarily neighbors, and the local objective of each node $i$ is a function of all the elements assigned to node $i$. In [34], each node’s local objective involves not only its own decision variable but also a common aggregative term associated with the decision variables of all the nodes. The global objective in [35] consists of a separable function and a known-to-all function whose variable is a linear combination of all nodes’ decision variables. The above-mentioned objective functions are different from that of problem (1), while problem (1) admits a more general constraint model than those of [32, 33, 34, 35]. In particular, [32, 34, 35] only consider unconstrained problems, [33] only allows for local constraints, yet (1) involves global constraints with coupled variables.
In fact, by letting each node maintain local copies of the decision variables of all the other nodes, problem (1) can be converted into a globally-coupled consensus optimization problem, so that a handful of existing algorithms are applicable (e.g.,[38, 39]). However, such a reformulation would significantly increase the costs in memory, computations, and communications [30], so we do not adopt this approach.
@@
Assumption 1 indicates that problem (1) is a smooth convex optimization problem with zero duality gap due to the Slater’s condition in Assumption 1(g). Note that we only assume convexity for the global objective function, so that each local objective $f_{i}$ is not necessarily convex. Also, the Lipschitz conditions of each $\nabla f_{i}$, $\nabla g_{ij}$, and $g_{i}$ in Assumption 1(c), 1(d), and 1(e) only need to hold on the compact constraint set $X_{\mathcal{N}_{i}}$. The conditions in Assumption 1 are mild and, later in Section IV-B, we will compare them with the assumptions imposed in the related works.

### v4

@@ Abstract
Many realistic decision-making problems in networked scenarios, such as formation control and collaborative task offloading, often involve complicatedly entangled local decisions, which, however, have not been sufficiently investigated yet. Motivated by this, we study a class of globally-constrained decentralized optimization problems with a variable coupling structure that is new to the literature. Specifically, we consider a network of nodes collaborating to minimize a global objective subject to a collection of global inequality and equality constraints, which are formed by the local objective and constraint functions of the nodes. On top of that, we allow such local functions to depend on not only the corresponding node’s decision variable but the decisions of its neighbors as well. To address this problem, we propose a decentralized projected primal-dual algorithm, which incorporates gradient projection and virtual-queue techniques with a primal-dual-primal scheme. Under mild conditions, we derive $O(1/k)$ convergence rates for both objective error and constraint violations. Finally, two numerical experiments corroborate our theoretical results and illustrate the competitive performance of the proposed algorithm.
@@ I Introduction
Decentralized optimization has garnered considerable recent attention due to its broad applications in various networked systems such as communication networks[Rabbat2004], smart grids[Magnusson2016], and computing networks[boyd2011distributed]. In these networked systems, each node possesses a collection of local data, typically inaccessible to other nodes for privacy protection, and all the nodes aim to make optimal decisions by solving a global optimization problem determined by all their local data. Decentralized optimization techniques allow the nodes to collaboratively solve such a problem through communicating with their neighboring nodes only, leading to high scalability with respect to the network size and the data volume.
@@
To date, a host of decentralized optimization algorithms have been developed. Among these works, the majority focus on consensus optimization, which aims at finding a common decision that minimizes the sum of the local objectives under, mostly, no constraints [Nedic2017, ShiW2015] or rather simple constraints such as set constraints [WuXY2019, mai2019distributed], local equality/inequality constraints[yang2016multi, khatana2022dc], and equality/inequality constraints known to all the nodes [zhu2011distributed, yuan2015regularized]. Another line of research explores problems with local variables subject to coupling equality/inequality constraints [Lakshmanan2008, Nedic2018, LinX2006, THChang2014, Falsone2017, Notarnicola2019, LiangS2019, falsone2023augmented, LiangS2019a, camisa2021distributed, LiuC2020, Mateos2016, wang2023distributed, wu2022distributed]. In particular, to solve the challenging problems with globally-coupled nonlinear inequality constraints consisting of all the local constraint functions from the nodes, the distributed primal-dual subgradient algorithms [THChang2014, Falsone2017, Mateos2016, LiuC2020, LiangS2019a, wang2023distributed, falsone2023augmented], the distributed method based on operator splitting[LiangS2019], the distributed integrated primal-dual proximal algorithm [wu2022distributed], and the relaxation and successive distributed decomposition method[Notarnicola2019] along with its extension [camisa2021distributed] are developed. These works also derive various convergence results, including asymptotic convergence [THChang2014, Falsone2017, Notarnicola2019, LiangS2019, falsone2023augmented] as well as $O(\ln k/\sqrt{k})$ [LiangS2019a], $O(1/\ln k)$ [camisa2021distributed], $O(1/\sqrt{k})$ [LiuC2020, Mateos2016, wang2023distributed], and $O(1/k)$ [wu2022distributed] convergence rates.
@@
A common limitation of the aforementioned decentralized optimization methods [Nedic2017, ShiW2015, WuXY2019, mai2019distributed, Lakshmanan2008, Nedic2018, yang2016multi, khatana2022dc, zhu2011distributed, yuan2015regularized, LinX2006, THChang2014, Falsone2017, Notarnicola2019, LiangS2019, falsone2023augmented, LiangS2019a, camisa2021distributed, LiuC2020, Mateos2016, wang2023distributed, wu2022distributed] is that each node’s local objective and constraint functions only rely on its own local decision variable. This excludes many realistic problems where each node’s local function values are influenced by the decisions made by other nodes in the network, especially the node’s neighbors. A representative example is the task offloading problem over cooperative mobile edge computing networks [xiao2018distributed], where each edge server’s task completion time and local resource consumption are affected by the task offloaded from the neighboring servers. Such a variable coupling structure also arises in collaborative resource allocation [huang2017collaborative], distributed model predictive control [WangCh2010], wireless localization [ShiW2010], etc.
A few existing works [santilli2018finite, todescato2020partition, pu2016quantization, HuJH2018, XiaoYY2017, lu2024distributed, rai2024distributed, srivastava2021network, ran2023distributed, wang2024distributed, borst2013nonconcave, Mota2014, Alghunaim2019, li2021distributed, du2025distributed, alghunaim2021dual] allow for various variable coupling structures in local functions. However, they only focus on unconstrained problems[santilli2018finite, todescato2020partition, li2021distributed], problems with local constraint sets[pu2016quantization], locally-coupled constrained problems[HuJH2018, XiaoYY2017, lu2024distributed, rai2024distributed, srivastava2021network, ran2023distributed, wang2024distributed, borst2013nonconcave, Mota2014, Alghunaim2019], or globally-coupled affine constraints [du2025distributed, alghunaim2021dual], where [XiaoYY2017, lu2024distributed, rai2024distributed, srivastava2021network, ran2023distributed] consider variable coupling in their local constraint functions but not in the local objectives. None of these works [santilli2018finite, todescato2020partition, pu2016quantization, HuJH2018, XiaoYY2017, lu2024distributed, rai2024distributed, srivastava2021network, ran2023distributed, wang2024distributed, borst2013nonconcave, Mota2014, Alghunaim2019, li2021distributed, du2025distributed, alghunaim2021dual] consider globally-coupled nonlinear constraints like those in [THChang2014, Falsone2017, Notarnicola2019, LiangS2019, falsone2023augmented, LiangS2019a, camisa2021distributed, LiuC2020, Mateos2016, wang2023distributed, wu2022distributed].
@@
Motivated by this, we consider a decentralized optimization problem with globally-coupled nonlinear inequality and linear equality constraints. Moreover, for each node, we allow both the local objective function and the local constraint functions to involve variable coupling, i.e., these local functions are associated with the decision variables of the node itself and all its neighbors. This coupling structure endows the optimization problem with a relatively general form that generalizes the prior prevalent decentralized optimization models in [Nedic2017, ShiW2015, WuXY2019, mai2019distributed, yang2016multi, khatana2022dc, zhu2011distributed, yuan2015regularized, Lakshmanan2008, Nedic2018, LinX2006, THChang2014, Falsone2017, Notarnicola2019, LiangS2019, falsone2023augmented, LiangS2019a, camisa2021distributed, LiuC2020, Mateos2016, wang2023distributed, wu2022distributed] with no variable coupling in the local functions and the variable coupling models in [santilli2018finite, todescato2020partition, pu2016quantization, HuJH2018, XiaoYY2017, lu2024distributed, rai2024distributed, srivastava2021network, ran2023distributed, wang2024distributed, borst2013nonconcave] with no constraints [santilli2018finite, todescato2020partition] or with only local constraints [pu2016quantization, HuJH2018, XiaoYY2017, lu2024distributed, rai2024distributed, srivastava2021network, ran2023distributed, wang2024distributed, borst2013nonconcave].
@@
In this paper, we develop a decentralized projected primal-dual algorithm for addressing the aforementioned globally-constrained optimization problem with variable coupling. To this end, we first construct a variant of the augmented Lagrangian function with respect to an equivalent form of the original problem, whose gradient has a certain degree of separability. Then, the primal update employs gradient projection to cause a descent of the augmented Lagrangian variant. The dual updates can be viewed as a non-trivial integration of ideas from the virtual queue method [YuHao2017] and a dual variant of the P-EXTRA algorithm [wu2019improved]. Eventually, the variable coupling structure and the globally-coupled constraints are effectively decoupled, so that our proposed algorithm can be executed in a fully decentralized fashion. We show that the proposed algorithm achieves $O(1/k)$ convergence rates in terms of both objective error and constraint violations under mild conditions. Its effectiveness and efficiency are demonstrated by two numerical examples.
@@
  • Compared to the existing works [THChang2014, Falsone2017, Notarnicola2019, LiangS2019, falsone2023augmented, LiangS2019a, camisa2021distributed, LiuC2020, Mateos2016, wang2023distributed, wu2022distributed] that are able to solve problems with global inequality constraints but no variable coupling (i.e., special cases of our problem), the $O(1/k)$ convergence rate of our proposed algorithm is stronger than the asymptotic convergence established in [THChang2014, Falsone2017, Notarnicola2019, LiangS2019, falsone2023augmented], faster than the convergence rates in [LiangS2019a, LiuC2020, camisa2021distributed, Mateos2016, wang2023distributed], and comparable to the rate in [wu2022distributed].
  • Our algorithm enjoys lower computational costs per iteration than many existing algorithms, because they require solving convex optimization problems at each iteration [wu2022distributed, wang2023distributed, LiangS2019a, falsone2023augmented, camisa2021distributed, Notarnicola2019, Falsone2017, LiuC2020], while our method only include simpler projection operations to update the primal variables. Additionally, our algorithm adopts constant step-sizes, which usually leads to faster convergence in practice compared to the algorithms with diminishing step-sizes [THChang2014, Falsone2017, Notarnicola2019, LiangS2019a, camisa2021distributed, LiuC2020, Mateos2016, wang2023distributed].
@@
Notation and Definition: For any set $X\subseteq\mathbb{R}^{d}$, $\operatorname{rel\;int}X$ is its relative interior. For any vectors $x,y\in\mathbb{R}^{n}$, we denote by $\max\{x,y\}$ the element-wise maximum of $x$ and $y$. Let $\mathcal{S}\subseteq\{1,\ldots,n\}$. Given $n$ column vectors $x_{1},\ldots,x_{n}$, $x_{\mathcal{S}}$ denotes the vector obtained by vertically stacking $x_{i}$ $\forall i\in\mathcal{S}$; given $n$ sets $X_{1},\ldots,X_{n}$, $X_{\mathcal{S}}$ denotes the Cartesian product of $X_{i}$ $\forall i\in\mathcal{S}$. We use $\|\cdot\|$ to denote the Euclidean norm and $\{\cdot,\cdot\}$ to represent an unordered pair. The projection of $x$ on a convex set $X$ is denoted by $\mathcal{P}_{X}(x)$, i.e., $\mathcal{P}_{X}(x):=\mathop{\arg\min}_{x^{\prime}\in X}\|x^{\prime}-x\|$. In addition, $I_{d}$ and $\mathbf{O}_{d}$ are the $d\times d$ identity matrix and all-zero matrix, respectively, and $\mathbf{1}_{d}$ ($\mathbf{0}_{d}$) is the $d$-dimensional all-one vector (all-zero vector), where the subscripts may be omitted for simplicity. For any real matrix $A\in\mathbb{R}^{m\times n}$, $A]_{ij}$ is its $(i,j)$-entry, $\operatorname{Range}(A)$ is its range, $\operatorname{Null}(A)$ is its null space, and $\|A\|$ is its spectral norm. If $A\in\mathbb{R}^{n\times n}$ is symmetric and positive semidefinite, $\|x\|_{A}:=\sqrt{x^{T}Ax}$ for any $x\in\mathbb{R}^{n}$.
@@ II Problem Formulation
For each $i\in\mathcal{V}$, $x_{i}\in\mathbb{R}^{d_{i}}$ is the decision variable of node $i$ and is restricted to node $i$’s constraint set $X_{i}\subseteq\mathbb{R}^{d_{i}}$. Additionally, $f_{i}:\mathbb{R}^{\sum_{j\in\mathcal{N}_{i}}d_{j}}\rightarrow\mathbb{R}$ is the local objective function and $g_{i}:\mathbb{R}^{\sum_{j\in\mathcal{N}_{i}}d_{j}}\rightarrow\mathbb{R}^{p}$ is the local inequality constraint function of node $i$, which are known only to node $i$ but are associated with the decision variables of both node $i$ and its neighbors, denoted by $x_{\mathcal{N}_{i}}$. Each $g_{i}$ is a vector-valued function, and we express it as $g_{i}\left(x_{\mathcal{N}_{i}}\right)\!=\![g_{i1}\left(x_{\mathcal{N}_{i}}\right),\ldots,g_{ip}\left(x_{\mathcal{N}_{i}}\right)]^{T}$. Moreover, $A_{i}\!\in\!\mathbb{R}^{m\times\sum_{j\in\mathcal{N}_{i}}d_{j}}$ and $b_{i}\in\mathbb{R}^{m}$, locally owned by node $i$, encode node $i$’s local component of the global equality constraint in (1), which, again, involves the decision variables of all the nodes in $\mathcal{N}_{i}$.
@@
Compared to most distributed optimization models[Nedic2017, ShiW2015, WuXY2019, mai2019distributed, yang2016multi, khatana2022dc, zhu2011distributed, yuan2015regularized, Lakshmanan2008, Nedic2018, LinX2006, THChang2014, Falsone2017, Notarnicola2019, LiangS2019, falsone2023augmented, LiangS2019a, camisa2021distributed, LiuC2020, Mateos2016, wang2023distributed, wu2022distributed, santilli2018finite, todescato2020partition, pu2016quantization, HuJH2018, XiaoYY2017, lu2024distributed, rai2024distributed, srivastava2021network, ran2023distributed, wang2024distributed, borst2013nonconcave] in the literature, problem (1) has a more complicated and general coupling structure and, thus, is more challenging to solve. Specifically, the problems considered in many existing works [Nedic2017, ShiW2015, WuXY2019, mai2019distributed, yang2016multi, khatana2022dc, zhu2011distributed, yuan2015regularized, Lakshmanan2008, Nedic2018, LinX2006, THChang2014, Falsone2017, Notarnicola2019, LiangS2019, falsone2023augmented, LiangS2019a, camisa2021distributed, LiuC2020, Mateos2016, wang2023distributed, wu2022distributed] are “separable” in the sense that each local function belonging to node $i$ is only associated with node $i$’s decision variable $x_{i}$. In contrast, problem (1) is non-separable, as each $f_{i}$, $g_{i}$, and $A_{i}$ is associated with $x_{\mathcal{N}_{i}}$, i.e., $x_{j}$ $\forall j\in\mathcal{N}_{i}$. In addition, $g_{i}$, $A_{i}$, $b_{i}$ $\forall i\in\mathcal{V}$ constitute the globally-coupled inequality and equality constraints in an additive manner. Such a non-separable structure is not uncommon in networked applications, as each node’s cost, utility, or capacity, etc., is often influenced by its neighbors’ states and plays a role in the network-wide problems. Similar variable coupling structures are also considered in [santilli2018finite, todescato2020partition, pu2016quantization, HuJH2018, XiaoYY2017, lu2024distributed, rai2024distributed, srivastava2021network, ran2023distributed, wang2024distributed, borst2013nonconcave], yet [santilli2018finite, todescato2020partition] only investigate unconstrained convex optimization, [pu2016quantization] focus on problems with local constraint sets, and [HuJH2018, XiaoYY2017, lu2024distributed, rai2024distributed, srivastava2021network, ran2023distributed, wang2024distributed, borst2013nonconcave] study locally-coupled constrained problems, where no variable coupling arises in the local objective functions in [XiaoYY2017, lu2024distributed, rai2024distributed, srivastava2021network, ran2023distributed].
@@
There have been several other coupling models for distributed convex optimization with an additive global objective function. In [Mota2014] and [Alghunaim2019], each element in a global decision vector is assigned to a subset of nodes that are not necessarily neighbors, and the local objective of each node $i$ is a function of all the elements assigned to node $i$. In [li2021distributed, du2025distributed], each node’s local objective involves not only its own decision variable but also a common aggregative term associated with the decision variables of all the nodes. The global objective in [alghunaim2021dual] consists of a separable function and a known-to-all function whose variable is a linear combination of all nodes’ decision variables. The above-mentioned objective functions are different from that of problem (1), while problem (1) admits a more general constraint model than those of [Mota2014, Alghunaim2019, li2021distributed, alghunaim2021dual, du2025distributed]. In particular, [Mota2014, Alghunaim2019] only allow for local constraints, [li2021distributed] only considers an unconstrained problem, [du2025distributed, alghunaim2021dual] only work for affine constraints, yet (1) involves both nonlinear inequality and linear equality global constraints with coupled variables.
In fact, by letting each node maintain local copies of the decision variables of all the other nodes or local copies of all the neighbors’ decision variables, problem (1) can be converted into a globally-coupled consensus optimization problem or an optimization problem with additional local consensus constraints, so that a handful of existing algorithms are applicable (e.g.,[li2020distributed, huang2023distributed, Alghunaim2019, wu2022distributed]). However, such reformulations would significantly increase the variable dimension and, thus, the costs in memory, computations, and communications [HuJH2018]. Therefore, we do not adopt such approaches.
@@
Assumption 1 indicates that problem (1) is a smooth convex optimization problem with zero duality gap due to the Slater’s condition in Assumption 1(g). Note that we only assume convexity for the global objective function, so that each local objective $f_{i}$ is not necessarily convex. Also, the Lipschitz conditions of each $\nabla f_{i}$, $\nabla g_{ij}$, and $g_{i}$ in Assumption 1(c), 1(d), and 1(e) only need to hold on the compact constraint set $X_{\mathcal{N}_{i}}$. The conditions in Assumption 1 are mild and commonly adopted in existing works[THChang2014, LiangS2019, wang2024distributed]. Many practical application problems satisfy the conditions in Assumption 1, including the two motivating examples in Section II-A. We will further compare these conditions with the assumptions adopted in related works in Section IV-B.